Subscribe to OSU Network Announcements feed
Announcements for OSU's Network and Computer Services
Updated: 10 min 50 sec ago

Milne Data Center Maintenance March 24th 6 to 7 AM

Thu, 03/19/2015 - 9:21am

Primary Contact: Joel Burks,

On March 24th at 6 AM the Network Operations team will be doing maintenance on two redundant switches in the Milne Data Center. The goal of this maintenance is to upgrade the firmware and implement best practices settings.

This maintenance will be performed with Cisco TAC online. During the maintenance, industry best practices will be followed to prevent outages on services that are dual-homed to the two network switches being upgraded. However, there may be a brief outage (up to 20 minutes) on devices that are only connected to a single network switch.

Emergency change for DNS, Exchange 3/10 @ 10PM

Tue, 03/10/2015 - 4:35pm

At 10pm tonight (March 10th) we will be moving two of OSU’s DNS servers and one Exchange client access server to a different data center. These servers will be taken down one at a time and will be unavailable for a couple of minutes each. The services are redundant and most customers will not notice any interruption in service.

This is an emergency change to restore redundancy to these critical services.

If you have any questions, please contact the OSU Computer Helpdesk.

-Kirsten Petersen
IT Manager, Information Services
Oregon State University

FormFusion Maintenance

Mon, 03/09/2015 - 8:14am

Primary Contact: Jill Swenson,

FormFusion will be unavailable from 5-6PM on Tuesday March 10 due to scheduled maintenance. We will be making configuration changes.

FormFusion users have been informed of this maintenace through the FormFusion users mail list.

OSU Wide Network Outage

Tue, 03/03/2015 - 10:06am

Notice of Unplanned Outage

Information Services experienced a failure in the network core during regularly scheduled maintenance this morning. We have restored some basic services at this time (9am) but are working on additional issues and still have many inter-dependent systems nonfunctional. We will notify the campus community when all services have been restored.


———————– Jonathan Dolan Director, IT Infrastructure

If you have questions or concerns about this outage, please contact Network Engineering at net (at) or call 737-HELP.

Network Outage 3/3

Tue, 03/03/2015 - 9:59am

Information Services suffered a catastrophic failure of the core network this morning while conducting regular announced maintenance starting at approximately 6:45am. The root cause of this failure was identified and services restored at approximately 9:25am. We are currently working with our vendor to understand the technical problems and will be conducting an architecture review and review of procedures to understand how this happened and how to avoid it in the future.

If you have any questions or concerns please call the OSU Computer Help Desk.


Jonathan Dolan | Director, IT Infrastructure | Information Services | Oregon State University
B211 Kerr Admin Building | Corvallis, OR 97331 | Phone: 541-737-5402

Banner Table Change

Wed, 02/25/2015 - 11:36am

Primary Contact: Jill Swenson,

On March 3 2015, between 9:00AM and 1:00PM, we will be redirecting a number of processes to point to a table which has been restructured (old table: SERSSGP, new table SYRSSGP). This change has been tested in DEVL. Updates include replacing one Banner INB form (SEAASGN will be replaced with SYASSGP) and updating various processes which currently use the old table.

The only impact visible to users will be the form change. The primary users of SEAASGN (Registrar’s office and ID Center) have been notified of the form change and have tested the new form in DEVL. The process changes will be verified, and the old table will remain in place until all dependencies are removed. If necessary, these changes can be easily reversed.

Milne Data Center Maintenance March 3rd 6 to 7 AM

Tue, 02/24/2015 - 11:17am

Primary Contact: Joel Burks,

We will be upgrading the Nexus 6000s in the Milne Data Center. Here is a list of the devices connected to the network gear.

SIG Eth1/4 c7k-mcc Bay 1 Port Eth1/6 #datadomain link1 Eth1/23 #mcc-b5-sw39.sig l Eth1/25 Po308 mlag1a-nic1. Eth1/26 Po308 mlag1b-nic1. Po #datadomain Po308 #mlag-domain1 Po309 #mcc-b5-sw39.sig o Eth105/1/10 #data domain appli Eth105/1/14 #GSA-mp01 external Eth105/1/15 #GSA-mp01 internal Eth110/1/3 mccA14-U06.sig.o.e Eth110/1/4 #ispvs-mp01.sig

ONID Eth1/44 #vm1-onid Eth110/1/10 gridc.ucsadm

SSG Eth100/1/1 ex3.oregonstate.ed Eth100/1/ #nws-sql3 Eth100/1/48 ex3.oregonstate.ed Eth105/1/2 ex3.oregonstate.ed

CN Eth100/1/4 #tcs1.scf Eth100/1/42 #tssdracsw1; z1-24 Eth105/1/3 #tss-exagrid.tss.o Eth105/1/4 #exagrid1a.tss.ore Eth105/1/5 #exagrid1b.tss.ore Eth105/1/6 #exagrid1c.tss.ore Eth105/1/7 #exagrid1d.tss.ore Eth105/1/18 #exagrid02.tss.ore

Facilities Eth100/1/13 #mobile-fm

NOC Eth100/1/16 aruba-3600-L6 Eth100/1/17 aruba-3600-L7 Eth100/1/18 aruba-3600-L8 Eth100/1/19 aruba-3600-L9 Eth100/1/ aruba-3600-L10 Eth100/1/31 aruba-3600-L11 Eth100/1/32 aruba-3600-L Eth105/1/16 #cp2.nws Eth105/1/37 #statseeker

Unknown Eth100/1/7 #TandbergContent Eth100/1/15 #MCC-2-Dell09-pd Eth100/1/27 #ares-thinclient; Eth100/1/33 IDC-SFTP.pos.orego Eth100/1/40 #tandberg-49a740 Eth105/1/9 #USB to_IP_applian

ONID Maintenance Monday February 23rd 11:00pm – 11:15pm

Mon, 02/23/2015 - 2:17pm

The ONID-FS Windows file service will be unavailable Monday February 23rd from 11:00pm to 11:15pm. The Samba service will be patched to fix a security vulnerability.

Andy Morgan ONID Support

Blackboard returned to service

Sat, 02/21/2015 - 8:17am

Scheduled Blackboard maintenance is complete, and services have been restored. If you have questions or concerns please contact

ONID Maintenance Saturday February 21st 8:00am – 8:30am

Fri, 02/20/2015 - 4:21pm

Maintenance will be performed on the following ONID services Saturday February 21st from 8:00am to 8:30am:

– ONID Webmail ( – MyONID website (ONID Password changes, ONID Sign-Up) – ONID SSO (the old SSO system, not CAS) – Cacti graphs (

An internal MySQL database server will be upgraded. These services will be unavailable while the MySQL upgrade is happening.

Andy Morgan ONID Support

Data Center Maintenance in B210 (Feb 24th 6 am to 7 am)

Wed, 02/18/2015 - 11:26am

Primary Contact: Joel Burks,

B210 Data Center will be undergoing maintenance between 6 and 7 am.

[DCA-All] Exchange All Clear

Mon, 02/16/2015 - 12:57am

Jason Appah Network Engineering 4200 Valley Library Oregon State University Corvallis, OR 97331 541-737-8801 [Description: cid:image001.png@01CD583B.8BD4B460]

The first client facing change for Exchange 2013 services was successful. There was a short disruption in service for 2-3 minutes starting at approximately 10:25PM.

We tested the following applications and platforms: iOS 8.1.3 – Success with Outlook App, Mail app.. Creating new and existing accounts Android Kit-Kat – Success with built in app auto-discover, outlook app existing account Android lollipop – Success with builtin app, (Gmail) w/ autodiscover Windows 7 – Success outlook 2013, OWA Server 2008r2 – outlook 2010, existing, OWA Windows 8.1 – Success outlook 2013, OWA OSX Maverick – Success with Built-In app (mail). Outlook 2015, Outlook 2011 Auto-discover and existing, OWA Windows Tablet 8.1/RT – success with built in mail. , OWA Pop3 – successful IMAP – successful

Jason Appah Network Engineering 4200 Valley Library Oregon State University Corvallis, OR 97331 541-737-8801 [Description: cid:image001.png@01CD583B.8BD4B460]

Data Center Maintenance in B210 (Feb 17th 6 am to 7 am)

Wed, 02/11/2015 - 2:52pm

Primary Contact: Joel Burks,

The data center in Kerr B210 will be undergoing maintenance on Feb 17th 6 to 7 am.

OSU_Secure emergency maintenance

Wed, 02/11/2015 - 7:20am

Service(s) affected: OSU_Secure wireless network.

Information Services conducted emergency maintenance this morning on OSU_Secure to address a failing authentication server. Mac and PC clients may warn of a security certificate change. This is expected behavior and does not represent a security risk.

If you have questions or concerns, please contact the OSU Computer Help Desk.

Outage Window: Start: 02/11/2015 6:00 AM End: 02/11/2015 6:15 AM

Reason for outage: Preventative maintenance.


Departments affected: All departments where the OSU_Secure wireless network is available.

Locations affected: All locations where the OSU_Secure wireless network is available.

Network Outage Policy

If you have questions or concerns about this outage, please contact Network Engineering at net (at) or call 737-HELP.

CWS Server Maintenance

Tue, 02/10/2015 - 12:33pm

Server maintenance for Central Web Services server is scheduled for 10:00pm on Wednesday, Feb 11. Affected websites are Drupal 6 sites and some non-Drupal CWS hosted sites and application. Drupal 7 sites will not be affected.

Affected databases are expected to be unavailable for less than 5 minutes.

Jean R Waters Assistant Director Academic Technology – Central Web Services Oregon State University 541-737-6588

2015-2-9 04:00-06:00

Thu, 02/05/2015 - 5:45pm

Primary Contact: Joel Burks,

NERO will be rebooting the La Grande Hub router on Monday, February 9th, between 0400-0600.

Customers connected to the La Grande Hub site will experience a brief loss of connectivity as the device resets.

La Grande Connected Customers: Eastern Oregon University Baker School District Baker County Public Library Blue Mountain Community College – Baker IMESD Baker City IMESD La Grande La Grande Public Library Moda/ODS Dental School – La Grande OSU Eastern Oregon Ag Research – Union OSU Baker Extension Service OSU Union Extension Service OSU Wallowa Extension Service

2015-2-5 04:00-06:00

Thu, 02/05/2015 - 9:55am

Primary Contact: Joel Burks,

NERO has learned of fiber splice work being performed on the fiber path between Eugene and Salem. Work is scheduled for Thursday, February 5th, between 0400-0600.

NERO does not expect this to be service impacting, but risk is elevated.

Data Center Maintenance in B210, A016 Kerr, and Milne

Tue, 02/03/2015 - 6:52am

Primary Contact: Joel Burks,

Changing MTU size on 4 links, turning off arp sync on router, and implementing new management vlan. No outage expected.